Solve this (600-800)/(600+800)(250-200)/(250+200) using the midpoint method
1
Expert's answer
2021-10-24T18:15:38-0400

Solution:

Using the mid-point method:

"=\frac{600 -800}{(600+800)/2 } \div \frac{250 -200}{(250+200)/2 }"


"=\frac{-200}{700 } \div \frac{50}{225 }"


"=\frac{-0.29}{0.22 } = -1.32"


= 1.32